OpenStudy (samara8954):

Jackson Middle School has 4 students in choir for every 1 student in band. If there are 46 students in band, how many students are in choir?

OpenStudy (samara8954):

2,116 184 50 42

OpenStudy (anonymous):

\[ \frac 14 = \frac{46}x \]

OpenStudy (samara8954):

help me answer

OpenStudy (anonymous):

1/4 = 46/x? For the answer you multiply 46*4 = 184

OpenStudy (samara8954):

Beth is making fruit salad. She adds 3 grapes for every 2 strawberries. If she uses 18 strawberries, how many grapes will she use?

OpenStudy (samara8954):

54 6 27 108

OpenStudy (anonymous):

3*18=54